A patient is having a magnetic resonance imaging scan (an MRI) and has neglected to remove a 8.1 cm diameter metallic bracelet. The resistance of the bracelet is 0.0190 Ω. The magnetic field within the MRI's solenoid is perpendicular to the opening of her bracelet. While taking the scan, the magnetic field decreases from 1.42 T to 0.37 T in 1.30 seconds. What is the magnitude of the average induced voltage in the bracelet?
_____???

Help please! Thanks!
Answer #1

Here,

radius of the bracelet, r = 8.1/2 = 4.05 cm

Now, for the induced current

induced current = change in magnetic flux/(time * resistance)

induced current = pi * 0.0405^2 * (1.42 - 0.37)/(1.30 * 0.0190)

induced current = 0.219 A

the induced current is 0.219 A